OpenStudy (anonymous):

help please on some questions? still learning find the quotient 10^-4/10^2

OpenStudy (anonymous):

10−4102 =10−4102 =10−4−2 =10−6 =1106 =11000000(Decimal: 0.000001)

OpenStudy (anonymous):

so awarer would be 1/1000000